# Identity Preserving 3D Head Stylization with Multiview Score Distillation

### v1

@@ Abstract
3D head stylization transforms realistic facial features into artistic representations, enhancing user engagement across applications such as gaming and virtual reality. While 3D-aware generators have made significant advancements, many 3D stylization methods primarily provide near-frontal views and struggle to preserve the unique identities of original subjects, often resulting in outputs that lack diversity and individuality. This paper addresses these challenges by leveraging the PanoHead model, synthesizing images from a comprehensive 360-degree perspective. We propose a novel framework that employs negative log-likelihood distillation (LD) to enhance identity preservation and improve stylization quality. By integrating multi-view grid score and mirror gradients within the 3D GAN architecture and introducing a score rank weighing technique, our approach achieves substantial qualitative and quantitative improvements. Our findings not only advance the state of 3D head stylization but also provide valuable insights into effective distillation processes between diffusion models and GANs, focusing on the critical issue of identity preservation. Please visit the project page.
@@ 1 Introduction
3D head stylization refers to transforming realistic facial and head features into artistic representations that can be rendered from multiple viewpoints, providing a dynamic and immersive experience. 3D head stylization has diverse applications in entertainment, advertising, education, enhancing user engagement, and creating relatable characters, and it is also a popular research topic due to its potential for innovation and impact across various fields.
Previously, face stylization was extensively explored in 2D image domains [56, 43, 13, 40], particularly through generative models like StyleGAN [24], which allowed for the manipulation of stylistic elements to create diverse, high-quality facial representations. With recent advancements in realistic 3D-aware face generators [8, 4], research in stylization has now shifted to the 3D domain [1, 25, 26, 61, 50, 5, 51, 29], as these models enable the creation of lifelike representations that can be viewed from multiple angles, significantly enhancing user experiences in applications such as gaming, virtual reality, and animation.
Most of these methods proposed for 3D stylization are built on the EG3D generator [8], which primarily synthesizes near-frontal views. This limitation restricts the generation of comprehensive 3D scenes from diverse viewpoints. In this work, we focus on PanoHead [4], which excels in synthesizing images from a 360-degree perspective. We establish applicable baselines using PanoHead, highlighting the challenges of stylizing 360-degree heads. Our findings demonstrate that most of the current methods fall short of effectively achieving comprehensive and cohesive stylization across all angles. This limitation arises because many of these methods rely on stylized images that are primarily available in near-frontal views that fine-tuned StyleGAN models generate [1, 50, 61].
There are other works that leverage text-based image diffusion or CLIP models to either generate datasets or guide training [25, 26, 5, 51, 29] providing more flexibility, especially for 3D generators. Especially, the Score Distillation Sampling (SDS) algorithm [45] has demonstrated its effectiveness in providing sufficient training feedback for the stylization of 3D portraits [51, 29]. It leverages the insights from text-based diffusion models to propagate signals through backpropagation. However, it has been noted that the SDS technique can result in a loss of diversity in the generated outputs [29]. For example, when the generator is fine-tuned for a Joker or Pixar style, the outputs from random samplings tend to exhibit significant similarity. Even though DiffusionGAN3D [29] introduces a relative distance loss to mitigate diversity loss, we find that its outputs, along with those from other competing methods, tend to be similar for different inputs. This results in challenges with identity preservation, as illustrated in Fig. 1. The outputs tend to be similar across different inputs, resulting in a loss of the unique identity of the original images.
@@
  • We propose using distillation with negative log-likelihood distillation (LD) [20] for domain adaptation of 3D-aware image generators, yielding sharper and more ID-preserving results compared to SDS [45].
  • We propose rank weighing for score tensors on latent channels to regularize the LD gradients, which we show achieve better input color and ID preservation in the distillation process.
@@
Our method has shown significant qualitative and quantitative improvements over the relevant head stylization methods and provides important insights into the distillation from diffusion to GAN backbones.
@@ 2 Related Work
3D-aware generators. Generative Adversarial Networks (GANs), when integrated with differentiable renderers, have made significant advancements in generating 3D-aware images consistent across multiple views. While earlier methods focus on mesh representations [41, 15, 11, 12], the latest innovations leverage implicit representations [9, 14, 39, 36]. Among implicit representations, triplane representations have become particularly popular due to their computational efficiency and the high quality of the generated outputs [8, 4]. The architectures of models like EG3D [8] and PanoHead [4] are reminiscent of the StyleGAN2 structure [24]. These frameworks consist of mapping and synthesis networks that create triplanes, which are then projected into 2D images through volumetric rendering processes similar to those used in NeRF [34]. While EG3D is trained on the FFHQ dataset [23] with limited angle diversity, PanoHead allows for a 360-degree perspective in face generation, attributed to its dataset selection and model enhancements. In our work, we investigate PanoHead to achieve 3D portrait stylization. We are interested in 3D portrait stylization of real input images, which requires embedding images into PanoHead’s latent space. This task is referred to as image inversion and extensively studied for StyleGAN generators [2, 46, 3, 42, 58, 59] and recently for 3D-aware GAN models [54, 60, 6, 30, 7].
@@
Adopting image generators to new domains. A popular method for stylizing 2D and 3D portraits involves fine-tuning generators initially trained to produce realistic faces. This process, known as domain adaptation of generators, utilizes the generators’ expertise in synthesizing real images. It can be accomplished using either a small number of samples for each style [31, 37] or by employing text-guided models to direct the training process [25, 26, 5, 51, 29]. Using text-based models to guide stylization offers the advantage of generating styles based on various prompts, which can be challenging to sample directly. These models, trained on extensive datasets, effectively acquire knowledge in a disentangled manner, enabling them to produce meaningful results for intriguing prompts that may not have corresponding data available [48, 44, 38]. Moreover, this guidance can be provided from multiple poses for 3D-aware generators, allowing for flexibility beyond the poses in the collected datasets. While previously, CLIP loss has been a popular choice for the guidance [35, 19, 49], more recently, the Score Distillation Sampling (SDS) algorithm has demonstrated impressive performance [45, 51, 29]. However, previous research indicates that when generators are fine-tuned solely using SDS losses, they may collapse into producing a fixed image pattern, irrespective of the input noise [51, 29]. To promote diversity, both StyleGANFusion [51] and DiffusionGAN3D [29] introduce regularizers aimed at enhancing variety. However, we observe that this diversity remains constrained, leading to issues with identity preservation, where different inputs yield similar outputs. In this work, we propose a framework that enables 3D portrait stylization guided by text prompts while maintaining identity preservation.
@@ 3 Method
Figure 2: Our proposed training approach with mirror gradients (a) and grid distillation (b). Dashed and non-dashed lines show backpropagation and forward paths. G, R, SR, and SD denote generator, neural renderer, super-resolver, and denoising UNet, respectively. Output of SD is the score, $\nabla_{x_{t}}\log p(x_{t}^{\pi}|y)$. We employ a depth-conditioned ControlNet in (b).
Our method involves fine-tuning the pre-trained parameters of PanoHead to adapt it for generating images from different domains. This section will go through the development of the method. Specifically, we will explain the adoption of LD to 3D-aware image generators, mention the differences between SDS, extend LD with cross-pose dependencies and grid denoising, and use rank reduction on score tensors.
@@ 3.1 Preliminaries
$\mathbb{E}_{x_{0},\epsilon,t}\{\|\epsilon-\hat{\epsilon}(x_{t},t)\|^{2}\}$ (2)
Score-based stochastic differential equations (SDE). In the score-based SDE framework [52], the forward process (Eq. 3) adds Gaussian noise to $x$ over time, where $g(t)=\sqrt{1-\bar{\alpha_{t}}}$ is the noise scale and $w$ is Wiener process. The reverse process (Eq. 4) is modeled as a reverse-time SDE and denoises $x$ by moving in the direction of higher probability density, where $\nabla_{x_{t}}\log p(x|y)$ is the score function. We keep the drift terms $f=0$ for simpler notation.
@@
$\nabla_{x_{t}}\log p(x|y)\approx\frac{-\hat{\epsilon}(x_{t},t|y)}{g(t)}$ (5)
@@ 3.2 ID-Preserving Stylization with Distillation
Likelihood distillation (LD) objective. We adopt a part of the distillation procedure in [20] explained in this section briefly. A detailed derivation is provided in Supplementary for completeness.
@@
where $\frac{\partial x_{t}^{\pi}}{\partial x_{0}^{\pi}}$ is $\sqrt{\bar{\alpha_{t}}}$ from Eq. 1. Notice that to update $\theta$, we do not need to back-propagate through the denoiser and can acknowledge the output as a part of the gradient. For the domain adaptation task, we assign $\theta$ as the PanoHead layers.
Note that LD is different than SDS [45] as it optimizes negative log-likelihood instead of reverse KL-divergence. Compliant with previous observations [53], we also observe that SDS yields blurry results, as LD is sharp in domain adaptation tasks. We attribute this to the fact that SDS is inherently mode-seeking due to the subtraction of ground truth noise $\epsilon$ from the estimated noise $\hat{\epsilon}$ and high classifier free guidance (CFG) [18] weight to avoid divergence [45]. In comparison, LD is diversity-seeking by not utilizing $\epsilon$, and does not require very high CFG weights, promoting diversity among samples.
We also note that even though LD yields sharper results, it is also said to suffer from stability in longer training regimes; hence, a balanced score distillation approach on top of Eq. 10 is proposed [20]. However, we observe that the strong prior of the GAN backbone is enough to avoid divergence with LD.
Rank weighted score tensors. While LD reduces the smoothness issues, we come across some artifacted results, especially around the hair, ears, and neck on some prompts (Fig. 3, full-rank). Notice that the artifacts are more focused on incorrect color distribution rather than the style itself. Inspired by previous work utilizing feature decomposition techniques [21, 27] and expecting a disentanglement, we investigate the SVD of the score tensors along the VQ-VAE latent channel (4) dimension. We notice that the first rank contains most of the stylization and conclude that surpassing the contributions of the lower three ranks can mitigate undesired tints. Hence, we re-weigh the diagonal singular value matrix S with linearly decaying coefficients W from the largest singular value to the smallest. Specifically,
@@
As seen in Fig. 3, rank-weighted score tensors eliminate the problematic tints and set a good baseline for the following improvements. We set the weight scores based on our empirical analysis, and the same weights are used for the training of all prompts.
@@
$+\sum_{\pi\neq\pi^{\prime}}\frac{\partial\log p(x_{t}^{\pi}|y)}{\partial x_{0}^{\pi^{\prime}}}\frac{\partial x_{0}^{\pi^{\prime}}}{\partial\theta}\}$
Notice that $\frac{\partial\log p(x_{t}^{\pi}|y)}{\partial x_{0}^{\pi^{\prime}}}=\nabla_{x_{t}}\log p(x_{t}^{\pi}|y)\frac{\partial x_{t}^{\pi}}{\partial x_{0}^{\pi^{\prime}}}$ in Eq. 12:
@@
By utilizing the symmetry prior for human heads, we assume that if $\pi$ and $\pi^{\prime}$ are yaw-symmetric camera matrices, $x_{t}^{\pi}=\mathbf{M}(x_{t}^{\pi^{\prime}})$ where $\mathbf{M}$ is the vertical mirror (flip) operator. Then, by Eq. 1, $\frac{\partial x_{t}^{\pi}}{\partial x_{0}^{\pi^{\prime}}}$ simply becomes $\mathbf{M}\sqrt{\bar{\alpha_{t}}}$. Further simplifying the expression Eq. 13 yields:
@@
As shown in Fig. 2 (a), this intuitively means that we utilize the same score estimation for mirror poses but must also mirror the gradients while back-propagating to the generator. Since 3D-consistent face domain editing with 2D-diffusion models is challenging, keeping the same score estimation for different renders of the same ID under logical constraints avoids deviating from the convergence path. This is because there is no guarantee that $\nabla_{x_{t}}\log p(x_{t}^{\pi}|y)$ and $\nabla_{x_{t}}\log p(x_{t}^{\pi^{\prime}}|y)$ will indicate the same direction for the same $y$.
@@
where $\{x_{0}^{\pi}\}$ is defined to have all poses $\{0,1,...,N\}$. However, this is not computationally feasible. Instead, as shown in Fig. 2 (b), we approximate $\{x_{0}^{\pi}\}$ with a 2$\times$2 grid, where each element is a $x_{0}^{\pi}$ with different render pose $\pi$. This way, denoising UNet can correlate between different renders of $\theta$, improving stylization consistency across views. We further employ a depth-conditioned ControlNet [62] as a regularizer to ensure the estimated score does not collapse to a single fused image.
@@
Effect of super-resolution network in style-based 3D GANs. The PanoHead model generates images at a resolution of $512\times 512$ [4], and the diffusion model similarly processes $512\times 512$ images when using LD-losses [47]. For grid construction, we arrange four images in a $2\times 2$ layout but reduce the individual image size to $256\times 256$ to avoid memory issues during forward passes. This way, the final grid image has the size of $512\times 512$. Although this maintains a good correlation between poses and consistent stylization, it creates a resolution mismatch when gradients are directly propagated from the PanoHead output. To resolve this, we experiment with feeding the gradients before the SR network, where the renderer outputs images at a lower resolution of $64\times 64$. This adjustment is only necessary for multi-view distillation, where we backpropagate $\nabla_{\theta}\mathcal{L}_{\text{LD}_{g}}$ before the SR layers. For mirror poses, we do not need to skip the super-resolution layers since they match the correct resolution, and so we backpropagate $\nabla_{\theta}\mathcal{L}_{\text{LD}}$ after the SR, as visualized in Fig. 2. This strategy improves the stylization quality and reduces unwanted artifacts, such as blur, over-saturation, and color tint.
@@
Figure 3: Ablation study on rank weighing. After SVD, four rank-1 matrices are obtained, and $\text{k}^{\text{th}}$ are chosen for reconstruction. Weighted rank improves the results compared to full rank. In this figure, we also present our final results for comparisons with LD + weighted rank and our multi-view and mirror gradient optimizations for completeness.
Figure 4: Qualitative stylization results provided in 360-degree views.

### v2

@@ Abstract
3D head stylization transforms realistic facial features into artistic representations, enhancing user engagement across applications such as gaming and virtual reality. While 3D-aware generators have made significant advancements, many 3D stylization methods primarily provide near-frontal views and struggle to preserve the unique identities of original subjects, often resulting in outputs that lack diversity and individuality. Leveraging the PanoHead model which provides 360-degree consistent renders, we propose a novel framework that employs negative log-likelihood distillation (LD) to enhance identity preservation and improve stylization quality. By integrating multi-view grid score and mirror gradients within the 3D GAN architecture and introducing a score rank weighing technique, our approach achieves substantial qualitative and quantitative improvements. Our findings not only advance the state of 3D head stylization but also provide valuable insights into effective distillation processes between diffusion models and GANs, focusing on the critical issue of identity preservation. Please visit the project page.
@@ 1 Introduction
3D head stylization refers to transforming realistic facial and head features into artistic representations that can be rendered from multiple viewpoints. 3D head stylization has diverse applications in entertainment, augmented-reality, advertising, education, and creating relatable characters, making it a popular research topic.
Previously, face stylization was extensively explored in 2D image domains [57, 14, 40], particularly through generative models like StyleGAN [25], which allowed for the manipulation of stylistic elements to create diverse, high-quality facial representations. With recent advancements in realistic 3D-aware face generators [9, 4], research in stylization has now shifted to the 3D domain [1, 26, 27, 62, 49, 5, 50, 29], as these models enable the creation of lifelike representations that can be viewed from multiple angles, significantly enhancing user experiences.
Most of these methods proposed for 3D stylization are built on the EG3D generator [9], which primarily synthesizes near-frontal views. This limitation restricts the generation of 3D scenes from diverse viewpoints. In this work, we focus on PanoHead [4], which excels in synthesizing images from a 360-degree perspective. We establish applicable baselines using PanoHead, highlighting the challenges of stylizing 360-degree heads. Our findings demonstrate that most current methods fall short of effectively achieving cohesive stylization across all angles. This limitation arises because many of these methods rely on stylized images that are primarily available in near-frontal views that fine-tuned StyleGAN models generate [1, 49, 62].
There are other works that leverage text-based image diffusion or CLIP models to either generate datasets or guide training [26, 27, 5, 50, 29] providing more flexibility, especially for 3D generators. Especially, Score Distillation Sampling (SDS) [44] has demonstrated its effectiveness in stylization of 3D portraits [50, 29], guiding the 3D representation to higher-density regions of the text-conditioned diffusion model. However, it has been noted that SDS can result in a loss of diversity in the generated outputs [29]. For example, when the generator is fine-tuned for a Joker or Pixar style, the outputs from random samplings tend to exhibit significant similarity. Even though DiffusionGAN3D [29] introduces a relative distance loss to mitigate diversity loss, we find that its outputs, along with those from other competing methods, tend to be similar for different inputs. This results in challenges with identity preservation, as illustrated in Fig. 1. The outputs tend to be similar across different inputs, resulting in a loss of the original identity.
@@
  • We propose using distillation with negative log-likelihood distillation (LD) [22] for domain adaptation of 3D-aware generators, yielding sharper and more ID-preserving results compared to SDS [44]. Improving ID preservation has not been done before on GANs via changing distillation characteristics; instead, the focus was diverted to additional losses to solve the underlying issues of SDS [50, 29].
  • We propose rank weighing for score tensors on latent VQ-VAE channels to regularize the LD gradients, which we show achieve better input color and ID preservation in the distillation process. This introduces a new approach to control stylization via distillation.
@@
Our method has shown significant qualitative and quantitative improvements over the relevant head stylization methods, presents good practices for distillation in one-step generators, and important insights into distillation from diffusion to GAN backbones.
@@ 2 Related Work
3D-aware generators. Generative Adversarial Networks (GANs), when integrated with differentiable renderers, have made significant advancements in generating 3D-aware images consistent across multiple views. While earlier methods focus on mesh representations [41, 17, 12, 13], the latest innovations leverage implicit representations [10, 15, 39, 36]. Among implicit representations, triplane representations have become particularly popular due to their computational efficiency and the high quality of the generated outputs, like EG3D [9] and PanoHead [4] which are reminiscent of the StyleGAN2 structure [25]. These frameworks consist of mapping and synthesis networks that create triplanes, which are then projected into 2D images through volumetric rendering processes similar to those used in NeRF [34]. While EG3D is trained on the FFHQ dataset [24] with limited angle diversity, PanoHead allows for a 360-degree perspective in face generation, attributed to its dataset selection and model enhancements. In our work, we investigate PanoHead to achieve 3D portrait stylization. We are interested in 3D portrait stylization of real, in-the-wild input images, which requires embedding images into PanoHead’s latent space. This task is referred to as image inversion and extensively studied for StyleGAN generators [2, 45, 3, 42, 59, 60] and recently for 3D-aware GAN models [55, 61, 6, 30, 7].
@@
Adopting image generators to new domains. A popular method for stylizing 2D and 3D portraits involves fine-tuning generators initially trained to produce realistic faces. This process, known as domain adaptation of generators, can be accomplished using either a small number of samples for each style [31, 37] or by employing text-guided models to direct the training process [26, 27, 5, 50, 29]. Using text-based models to guide stylization offers the advantage of generating styles based on various prompts, which can be challenging to sample directly. These models, trained on extensive datasets, effectively acquire knowledge in a disentangled manner, enabling them to produce meaningful results for intriguing prompts that may not have corresponding data available [47, 43, 38]. Moreover, this guidance can be provided from multiple poses for 3D-aware generators, allowing for flexibility beyond the poses in the collected datasets. While previously, CLIP loss has been a popular choice for the guidance [35, 21, 48], more recently, the Score Distillation Sampling (SDS) has demonstrated impressive performance [44, 50, 29]. However, previous research indicates that when generators are fine-tuned solely using SDS, they may collapse into producing a fixed image pattern, irrespective of the input noise [50, 29]. To promote diversity, both StyleGANFusion [50] and DiffusionGAN3D [29] introduce regularizers aimed at enhancing variety. However, we observe that this diversity remains constrained, leading to issues with identity preservation, where different inputs yield similar outputs. In this work, we propose a framework that enables 3D portrait stylization guided by text prompts while maintaining identity preservation.
@@ 3 Method
Figure 2: Our proposed training approach with mirror gradients (a) and grid distillation (b). Dashed and non-dashed lines show backpropagation and forward paths. G, R, SR, and SD denote generator, neural renderer, super-resolver, and denoising UNet, respectively. Output of SD is the score, $\nabla_{x_{t}}\log p(x_{t}^{\pi}|y)$. We employ a depth-conditioned ControlNet in (b). For the sake of simplicity, we omit CFG here.
Our method involves efficient fine-tuning the pre-trained parameters of PanoHead for generating images from different domains. This section will go through the development of the method. Specifically, we will explain the adoption of LD to 3D-aware image generators, mention the differences between SDS, extend LD with mirror pose dependencies and grid denoising, and use rank weighing on score tensors.
@@ 3.1 Preliminaries
$\mathbb{E}_{x_{0},\epsilon,t}[\ \|\epsilon-\hat{\epsilon}(x_{t},t)\|^{2}]$ (2)
Score-based stochastic differential equations (SDE). In the score-based SDE framework [51], the forward process Eq. 3 adds Gaussian noise to $x$ over time, where $g(t)=\sqrt{1-\bar{\alpha_{t}}}$ is the noise scale and $w$ is Wiener process. The reverse process, Eq. 4, is modeled as a reverse-time SDE and denoises $x$ by moving in the direction of higher probability density, where $\nabla_{x_{t}}\log p(x|y)$ is the score function and $y$ is the condition signal (text prompt). We keep the drift terms $f=0$ for simpler notation.
@@
$\nabla_{x_{t}}\log p(x|y)\approx{-{g(t)}^{-1}\hat{\epsilon}(x_{t};t,y)}$ (5)
@@
Score distillation sampling (SDS). SDS [44] leverages pre-trained 2D diffusion models to guide the generation of content in other domains, such as 3D, shown in Eq. 6:
$\nabla_{\theta}L_{\text{SDS}}(\phi,x=g(\theta))\triangleq\mathbb{E}_{t,\epsilon}\big{[}\omega(t)(\hat{\epsilon}(x_{t};y,t)-\epsilon)\frac{\partial x}{\partial\theta}\big{]}$ (6)
where $\omega(t)$ weights the timestep contribution, $\theta$ is the 3D representation, and $g$ is the differentiable renderer to project 3D to 2D. Intuitively, SDS guides $\theta$ toward higher-density regions in the diffusion model’s learned distribution.
@@ 3.2 ID-Preserving Stylization with Distillation
Likelihood distillation (LD) objective. We partially adapt the distillation procedure in [22] and tailor it to GANs, explained in this section. A detailed derivation is provided in Supplementary for completeness.
@@
where $\frac{\partial x_{t}^{\pi}}{\partial x_{0}^{\pi}}$ is $\sqrt{\bar{\alpha_{t}}}$ from Eq. 1. Notice that to update $\theta$, we do not need to back-propagate through the denoiser and can acknowledge the output as a part of the gradient. For this task, we assign $\theta$ as the PanoHead layers.
SDS vs LD on GANs. Note that LD is different than SDS [44] as it optimizes negative log-likelihood instead of reverse KL-divergence. Compliant with previous observations [54], we also observe that SDS yields blurry results, as LD is sharp in domain adaptation tasks. We attribute this to the fact that SDS is inherently mode-seeking due to the subtraction of ground truth noise $\epsilon$ from the estimated noise $\hat{\epsilon}$ and high classifier free guidance (CFG) [20] weight to avoid divergence [44]. In contrast, LD is diversity-seeking (by not utilizing $\epsilon$), does not require very high CFG weights, and is more suitable for our task considering that we utilize a GAN prior. This issue had not been investigated before on GANs, and the focus was diverted to additional losses to solve the underlying issues of SDS, e.g. [50, 29].
We also note that [22] proposes multiple-gradient descent on top of Eq. 11, which we simply omit since the pretrained GAN backbone is strong enough to avoid divergence. However, for the sake of completeness, we also compare it with LD in Fig. 3, along with our full pipeline. In our comparisons, LD and PlacidDreamer alter expressions (e.g., opening a closed mouth in Fig. 3, row 1, or removing visible teeth, row 2), misrepresent accessories like glasses, and distort features such as hair in stylized domains. Our full pipe preserves these details more accurately.
Input LD PlacidDreamer Ours
Wolf
Joker
Pixar
Figure 3: Difference between LD, PlacidDreamer’s additions on LD, and our additions on LD.
Rank weighted score tensors. While LD reduces the smoothness issues, we may come across some artifacted results, especially around the hair, ears, and neck on some prompts (Fig. 5, full-rank). Notice that the artifacts are more focused on incorrect color distribution rather than the style itself. Since diffusion latents must also contain color distribution information [52], we decide to investigate the SVD of the score tensors along the VQ-VAE latent channel (4) dimension. We notice that the first rank contains most of the stylization and surpassing the contributions of the lower three ranks can mitigate undesired tints. Hence, we re-weigh the diagonal singular value matrix S with linearly decaying coefficients W from the largest singular value to the smallest. Specifically,
@@
As seen in Fig. 5, rank-weighted scores eliminate the problematic tints and set a good baseline for the following improvements. We set the weight scores based on our empirical analysis, and use the same weights for all prompts. We also show that our weighing generalizes to other latent spaces in Fig. 4, revealing that emphasizing the largest rank-1 matrix via weighted W yields more consistent stylization. More results are in Supplementary.
Input SD1.5 RV3.0 RV5.1 SDXL
Pixar
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Statue
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Distillation from different latent spaces, with default W and (left) weighted $\mathbf{W}=\text{diag}(1,0.75,0.5,0.25)$ (right).
@@
By utilizing the symmetry prior for human heads, we assume that if $\pi$ and $\pi^{\prime}$ are yaw-symmetric camera matrices, then $x_{t}^{\pi}=\mathbf{M}(x_{t}^{\pi^{\prime}})$ where $\mathbf{M}$ is the vertical mirror (flip) operator. Then, by Eq. 1, $\frac{\partial x_{t}^{\pi}}{\partial x_{0}^{\pi^{\prime}}}$ simply becomes $\mathbf{M}\sqrt{\bar{\alpha_{t}}}$. Further simplifying the expression Eq. 13 yields:
@@
As shown in Fig. 2 (a), this intuitively means that we utilize the same score estimation for mirror poses but must also mirror the gradients while back-propagating to the generator. Since 3D-consistent face domain editing with 2D-diffusion models is challenging, keeping the same score for different renders of the same ID under logical constraints avoids deviating from the convergence path. This is because $\nabla_{x_{t}}\log p(x_{t}^{\pi}|y)$ and $\nabla_{x_{t}}\log p(x_{t}^{\pi^{\prime}}|y)$ indicating the same direction for the same $y$ is not guaranteed.
@@
where $\{x_{0}^{\pi}\}$ is defined to have all poses $\{0,1,...,N\}$. However, this is not computationally feasible. Instead, as shown in Fig. 2 (b), we approximate $\{x_{0}^{\pi}\}$ with a 2$\times$2 grid, where each element is a $x_{0}^{\pi}$ with different render pose $\pi$. This way, denoising UNet can correlate between different renders of $\theta$, improving stylization consistency across views. We further employ a depth-conditioned ControlNet [63] to ensure the estimated score does not collapse to a single fused image. We also leverage CFG [20] on each backpropagation path in Fig. 2, but omit from figure for simplicity.
@@
Effect of super-resolution network in style-based 3D GANs. The PanoHead model generates images at a resolution of $512\times 512$ [4], and the diffusion model similarly processes $512\times 512$ images when using LD [46]. For grid construction, we arrange four images in a $2\times 2$ layout but reduce the individual image size to $256\times 256$ to avoid memory issues during forward passes. This way, the final grid image has the size of $512\times 512$. Although this maintains a good correlation between poses and consistent stylization, it creates a resolution mismatch when gradients are directly propagated from the PanoHead output. To resolve this, we experiment with feeding the gradients before the SR network, where the renderer outputs images at a lower resolution of $64\times 64$. This adjustment is only necessary for multi-view distillation, where we backpropagate $\nabla_{\theta}\mathcal{L}_{\text{LD}_{g}}$ before the SR layers. For mirror poses, we do not need to skip the super-resolution layers since they match the correct resolution, as visualized in Fig. 2. This strategy improves the stylization quality and reduces unwanted artifacts, such as blur, over-saturation, and color tint, seen in Fig. 10.
@@
Figure 5: Ablation study on rank weighing. After SVD, four rank-1 matrices are obtained, and $\text{k}^{\text{th}}$ are chosen for reconstruction. Weighted rank improves the results compared to full rank. We also present final results (ours) for comparisons with LD + weighted rank and our multi-view and mirror gradient.
Figure 6: Qualitative stylization results of 3D domain-adaptation methods, provided in 360-degree views.